WACC and Percentage of Debt Financing Hook Industries' capital structure consists solely of debt and common equity. It can issue debt at rd = 8%, and its common stock currently pays a $3.50 dividend per share (D0 = $3.50). The stock's price is currently $35.00, its dividend is expected to grow at a constant rate of 8% per year, its tax rate is 35%, and its WACC is 12.30%. What percentage of the company's capital structure consists of debt? Round your answer to two decimal places.
Answer #1

After tax cost of debt = rd * ( 1-tax rate)

= 8% * (1 - 35%)

= 5.20%

Cost of Equity = [Expected Dividend / Current Price] + Growth Rate

= [ $ 3.5 * (1.08) / $ 35] + 8%

= 18.80%

Let the weight of equity be x

Hence weight of debt = ( 1-x)

WACC = (Cost of Debt * Weight of Debt) + (Cost of Equity * Weight of Equity)

12.30 % = [ 5.20% * ( 1-x) ]+ (18.80% * x)

12.30 % = [5.20% - 5.20% x + 18.80% * x]

12.30% - 5.20% = (- 5.20% x + 18.80% * x)

x= 7.10/ 13.6

= 0.522058823

Debt = ( 1-x)

= 1- 0.522058823%

= 47.79%

Hence the correct answer is 47.79%